# D4C Glove-train: Solving the RPM and Bongard-logo Problem by Circumscribing and Building Distribution for Concepts

### v9

@@ VI-A D2C on Bongard-Logo problem
In the Bongard-Logo problem, the determination of image attributes is influenced by the specific contextual concepts associated with each individual case [32]. Distinct concepts require the network to encode specific pixel configurations into differentiated attribute representations, potentially compelling the network to maintain mutually exclusive representation encoding strategies and methods [32]. This scenario often leads to concept confusion within the network, rendering it uncertain about which concept to rely on when processing a Bongard-Logo case [32]. In a nutshell, from the perspective of network, there exists mutual infringement and erroneous coupling among the high-dimensional human concepts in the Bongard-Logo. Therefore, this study aims to develop a methodology that effectively decouples the concepts across different cases within the Bongard-Logo problem. We posit that by successfully decoupling such that distinct cases rely on distinct concepts we may alleviate, to a certain extent, the issue of the network losing concept matching when confronted with a Bongard-Logo problem.
To achieve this objective, we have redefined the concepts within the Bongard-Logo database, representing them as conditional probability distributions $p_{i}(x|\,Y_{i})$ and $q_{i}(x|\,Y_{i})$. Here, $i$ ranges from $1$ to $n$, denoting each index of the Bongard-Logo cases. Additionally, we introduce the notation $\{Y^{\prime}_{i}|\,i\in[1,n]\}$ to represent the set of reorganized concepts. It is evident that the number of reorganized concepts corresponds to the number of Bongard-Logo cases. Through this concept reorganization, we strive for mutually exclusive expressions of concepts across all Bongard-Logo cases. In other words, each case is categorized and solved based on its distinct concepts. By establishing a one-to-one correspondence between problem cases and these redefined concepts, we aim to enable the network to avoid uncertainty in determining which concept to rely on when solving a specific case. This approach aims to mitigate the impact of concept confusion on the solution for the Bongard-Logo problem.
@@ VI-A1 D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
By incorporating an additional ancestral sampling procedure into the traditional mini-batch training sampling methods, this study constructs a Bongard-logo problem with reorganized concept space for training purposes. This constitutes a hot-swappable approach for training the Bongard-Logo solver, which is designated as phase 1 of the D2C method in this paper.
Specifically, we denote a batch of Bongard-logo cases as $\{x_{ij}|\,i\in[1,b],j\in[1,14]\}$, where $i$ denotes the index of the Bongard-logo case within a batch, the batch size is $b$, and $j$ represents the sample index within a single Bongard-logo case, with $j\in[1,7]$ being the primary samples and $j\in[8,14]$ being the auxiliary samples. From this batch of cases $\{x_{ij}\}$, we deliberately choose two batches of cases $\{x_{tj}|\,j\in[1,14]\}$ and $\{x_{\tilde{t}j}|\,j\in[1,14]\}$, ensuring that $t$, $\tilde{t}$, and $i$ are all distinct. This process can be viewed as ancestral sampling. The extracted primary samples $\{x_{tj}|\,j\in[1,7]\}$ and $\{x_{\tilde{t}j}|\,j\in[1,7]\}$ are then incorporated into $\{x_{ij}\}$ as auxiliary samples, yielding an expanded set $\{x_{i\tilde{j}}\}$ where $\tilde{j}\in[1,28]$. Subsequently, The new sample groups $\{x_{i\tilde{j}}\}$, which oppose the concepts from different cases compared to their original form, can be fed into any discriminator used to solve Bongard-Logo problems, such as Triple-CFN[32] and PMoC[31]. The feedforward process, as illustrated in the accompanying diagram, involves a strategic misalignment of Bongard-logo problem batches to facilitate ancestral sampling during implementation. The alignment of the D2C phase 1 process with the baseline in DC is illustrated in figure 13.
@@
It is worth emphasizing that the improvement in model performance observed through the aforementioned approach is not solely attributed to an increase in the number of negative examples within each contrastive group. This paper posits that the incorporation of negative examples, which assist in delineating concept boundaries, holds the key to enhancing a model’s proficiency in abstract reasoning tasks. Merely augmenting the dataset with insignificant negative examples does not lead to substantial gains in model performance. In an attempt to enrich the dataset, this study generated a substantial amount of Bongard-logo virtual data using DDPM and integrated these as auxiliary data points within the original problems. However, subsequent analysis revealed that this methodology did not tangibly enhance model accuracy.
In summary, the phase 1 of the D2C approach embodies a straightforward yet effective concept: it introduces positive examples from other cases as negative samples in the current case. This strategy sends a clear and additional signal to Bongard-Logo discriminators that distinct problem cases should rely on differing concepts. As a result, the first phase of D2C is widely applicable to various Bongard-Logo discriminators, including Triple-CFN and PMoC.
@@ VI-A2 D2C phase 2: The concept delineation method based on a new loss function term
This paper posits that, in addition to altering the manner of sampling training data, establishing an extra loss function to constrain the similarity between the representations of the primary group samples across different cases within a batch of Bongard-Logo training data can implicitly achieve the fundamental objective of this paper.
Specifically, in this paper, we exclusively retain the representation of primary group within the set $\{z_{ij}|\,i\in[1,b],j\in[1,14]\}$, which specifically correspond to $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$. Subsequently, we randomly partition the primary representation of each case within $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$ into two distinct subsets: $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$, where $\{z_{ih}|\,i\in[1,b]\}\cup\{z_{i\tilde{h}}|\,i\in[1,b]\}=\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$ and $\{z_{ih}\}\cap\{{z}_{i\tilde{h}}\}=\emptyset$. By taking the mean of the vectors within $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$, we derive the corresponding vectors $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$. Ultimately, we compute InfoNCE loss function for constraining the similarity among $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$ as the additional loss function item. The specific calculation process can be explained by:
@@
$=-\sum^{b}_{i=1}\log\frac{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}}}{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}+\sum\nolimits_{\tilde{i}=1,\tilde{i}\neq i}^{b}{{e^{({v_{i}}\cdot{v_{\tilde{i}}})/t}}}}}$
@@
$\text{and}\,\,\{z_{ih}\}\cup\{z_{i\tilde{h}}\}=\{z_{ij}|\,j\in[1,7]\}$ (12)
This loss term is to decrease the similarity of primary group representations across different cases. The schematic representation of the feedforward procedure is provided in the figure 14.
Figure 14: Feedforward process of D2C phase 2
In essence, the phase 2 of D2C attempts to achieve a similar goal as its first phase by formulating a new loss function, approaching the task from a different angle. Nevertheless, this strategy renders the phase 2 of D2C primarily compatible with backbone that can be abstracted as $f(z|x)$. This adaptability enables a majority of image encoders, such as ResNet, ViT, and Triple-CFN[32], to be compatible with the complete D2C methodology. However, the application of PMoC[31], which relies on probability-based measures to assess image identity, is restricted to the first phase of D2C.
@@ VI-B D2C on RPM probelm
Furthermore, the spirit of D2C method can be integrated with Lico-Net by introducing images from other case statements as negative examples within the cases. However, the application of D2C in RPM discriminators is limited to its phase 1, as it becomes challenging to forcibly apply the phase 2 to these discriminators. This process is illustrated in figure 15.
@@
In summary, D2C is a straightforward approach. In the context of Bongard-Logo, where the determination of image attributes heavily relies on the concepts outlined by the primary and auxiliary groups within each case, D2C addresses this challenge by decoupling the concepts across different cases. Nevertheless, regarding the RPM problem, which emphasizes consistency in image attribute identification, this study maintains a pessimistic outlook on the efficacy of D2C in tackling RPM. In a nutshell, D2C is suitable for addressing the Bongard-Logo problem, where there exists mutual infringement and erroneous coupling of concepts across cases, but it is challenging to apply D2C to the RPM problem, which involves sharing a common set of concepts across cases. It is evident that D2C employs a precautionary approach, aiming to minimize mutual infringement and erroneous coupling among concepts, even at the cost of potential overcorrection. Finally, the D2C and D3C methodologies are not mutually exclusive. Specifically, D2C endeavors to redefine concept boundaries for reasoning problems. According to D3C, representing concepts in the form of distributions is a more plausible approach compared to utilizing vectors.
@@ VII Defense-Driven Discrimination of Distributed Concepts (D4C)
This paper aims to delineate more appropriate concept boundaries for the RPM and Bongard-Logo problems.
@@ VII-A Negative samples are not merely intended to confound the discriminator
This paper posits that it is worthwhile to consider the significance of positive and negative examples in abstract reasoning problems. Suppose that the correct solutions to each case in a reasoning problem follow a Gaussian distribution. Given that some reasoning problems may have multiple correct solutions, it is also possible that the distribution is a mixed Gaussian distribution. This is exemplified in datasets such as OIG and D9 from Raven database. One of the core challenges in RPM problems lies in the existence of multiple correct answers. While CRAB has eliminated this core difficulty through data cleaning in the RAVEN dataset, it has achieved noteworthy results on the cleaned version.
Currently, existing discriminative models for solving RPM and Bongard-Logo problems often adhere to a pattern where positive examples yield high probabilities when substituted into the case’s problem statement, while negative examples yield low probabilities. In a nutshell, they are models with probabilistic labeling scheme. We can view this pattern in the following way: the discriminator constructs a solution distribution for the cases presented in abstract reasoning problems. Within this distribution, positive examples are assigned higher probabilities, whereas negative examples are assigned lower ones. The endeavor to train the discriminator can be conceptualized as a process of optimizing and fine-tuning the solution distribution, with the ultimate objective of aligning it more closely with the underlying distribution of correct solutions for the given cases. Using positive and negative examples to estimate a Gaussian distribution through the cross-entropy loss function can be seen as leveraging them to optimize the following function:
@@
Where $x$ represents the positive or negative samples, $\mu$ and $\sigma$ represent the mean and standard deviation of the Gaussian distribution separately. It can be observed that when positive samples are substituted into the formula, we aim for the smallest possible result, which corresponds to minimizing the mean squared loss between sample $x$ and the mean of the correct solution $\mu$. Conversely, when negative samples are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma$ of the respective segment.
Therefore, this paper posits that in reasoning problems, the establishment of positive examples aids deep models in estimating the mean of the correct solution distribution, while negative examples serve to estimate the variance. The variance represents the distance between the boundaries of the distribution and its center. Correspondingly, we posit that negative examples play a crucial role in abstract reasoning tasks, serving to delimit the concepts presented in each case. This underscores the significance of establishing both positive and negative examples in abstract reasoning problems. This significance reflects a despairing fact that the upper limit of the discriminator’s reasoning accuracy in this abstract reasoning problem is determined by the quality of positive and negative examples in the question.
@@ VII-B High-quality negative samples lead to more rigorous concept boundaries
The D2C methodology proposed in this study has established a novel problem boundary for deep networks by means of resampling negative samples and incorporating a fresh loss function. Experimental findings tentatively indicate that this boundary is better suited for scenarios characterized by the absence of concept supervision signals, as compared to the boundaries delineated by the original negative samples in Bongard-Logo. Nonetheless, the precise tightness of the new boundary remains indeterminate.
In essence, D2C introduces a new variance to the solution distribution of the network, the magnitude of which relative to the variance of the correct solution distribution is indeterminate. In the context of the Bongard-Logo challenge, the variance provided by D2C appears closer to that of the correct solution distribution in the absence of concept supervision signals, as compared to the variance described by the original negative samples.
Evidently, the efficacy of D2C is contingent upon specific dataset attributes. As anticipated, empirical evidence suggests that this approach may be insufficient for delineating precise boundaries within the framework of the RPM problem. Contrary to the Bongard-Logo paradigm, RPM problems postulate a shared set of core concepts across distinct cases, ensuring discreteness and stability. This shared nature precludes the utility of methods like D2C in addressing RPM challenges. Specifically, image properties within RPM cases are rigidly governed by definitions of “rules” or “progressive patterns,” maintaining consistent attributions across all cases.
The Bongard-Logo problem exhibits varying identifications and determinations of image attribute due to differing clustering bases among cases, thereby augmenting problem complexity and difficulty. Incorporating positive samples from other cases as negative examples can, to some extent, delineate concept boundaries between distinct cases, adhering to a cautious approach aimed at minimizing concept infringement. However, this strategy proves inapplicable in addressing challenges presented by RAVEN and PGM due to substantial disparities across multiple dimensions.
In summary, the success of D2C in addressing the Bongard-Logo problem can be attributed to its indirect enhancement of the quality of negative samples within Bongard-Logo cases. Conversely, its limitations in tackling the RPM problem can also be ascribed to its insignificant improvement in the quality of negative examples within RPM scenarios. To address the limitations inherent in the D2C approach, this paper introduces the D4C (Defense-Driven Discrimination of Distributed Concepts) method. The underlying idea of D4C is to generate higher-quality negative samples and samples within abstract reasoning problems. This approach aims to delineate a more appropriate problem boundary for the discriminator as compared to the original problem setup in abstract reasoning. By leveraging novel mechanisms and strategies, D4C aims to deliver more precise and reliable definitions of concept boundaries and foster further advancements in the realm of abstract reasoning.
@@ VII-C Impelmention of D4C method
Effective Bongard-Logo discriminators abound, with our preceding works, such as PMoC and Triple-CFN, achieving notable successes. These discriminators seamlessly integrate with the D2C and D3C methodologies proposed in this study, and pertinent experiments have been conducted to explore their combinations. These discriminators excel in accurately identifying and classifying Bongard-Logo cases, significantly contributing to the advancement of research in this domain.
Regarding discriminators capable of handling adversarial tasks, the Wasserstein GAN, a prominent adversarial generative network, poses a noteworthy requirement. The discriminator must undertake the task of measuring the distance between samples generated by the generator and real samples. This necessitates that the discriminator satisfies Lipschitz continuity to a certain extent.
@@
Specifically, we employ a standard Transformer encoder to encode a Gaussian distribution based on the representations of the primary samples in a Bongard-Logo case, along with an optimizable vector. This encoding approach bears similarity to D3C, wherein we extract the optimizable vector from the attention results and process it through a linear feed-forward layer to obtain the mean and logarithm of variance for the Gaussian distribution. Subsequently, we sample multiple representations from this distribution as confusing representations to challenge the discriminator, which must exert maximum effort to counter them. The generator’s training objective is to generate representations for Bongard-Logo cases that achieve higher scores in the discriminator compared to the representations of the original negative examples in the case. Conversely, the discriminator’s training objective is unambiguous: to classify these generated representations as negative examples. The aforementioned approach can be seamlessly applied to the PMoC model, providing a direct and effective means of enhancing its performance. The alignment between the generator and PMoC is illustrated in Figure 16.
@@
In this figure, the “$f_{\theta}(z_{ij}|x_{ij})$” and “$g_{\omega}(p^{\prime}_{i}(z_{ij}|y)|\{z_{ij}\}_{j=1}^{6},z_{ij^{\prime}})$” represent the perceptron and logic calculation modules in PMoC, respectively[31].
For the generator in Triple-CFN, it is essential to decay the generated representations. Specifically, when the generator with the aforementioned structure produces a confusing representation for Triple-CFN, this confusing representation must be weighted and summed with the representations in the auxiliary group. After taking the average of the representations in the auxiliary group, it is weighted and summed with the confusing representation in a ratio of $0.9:0.1$ to decay the confusing representation. This decay process is necessary due to the mechanism of Triple-CFN, which defines the identity of an image based on the direction of its representation. This mechanism makes it easier for a generator based on the Transformer encoder to generate realistic confusing representations, potentially causing Triple-CFN to collapse. Therefore, we have designed the aforementioned decay process. The alignment between the generator and Triple-CFN is illustrated in Figure 17. In this figure, “$f_{\theta}(z_{ij}|x_{ij})$” represents the encoder in Triple-CFN.
@@
However, Lico-Net will encounter challenges posed by the generator, which produces confusing representations aimed at completing the $3\times 3$ reasoning matrix. To address these challenges, Lico-Net will undertake an assessment of the internal consistency within the progressive patterns of the $3\times 3$ reasoning matrix, which are completed by these confusing representations. Any inconsistencies or irregularities identified within these representations will be flagged as negative samples, indicating incorrect responses. Through this rigorous process, Lico-Net strives to accurately identify and reject perplexing representations, thereby augmenting its overall precision and reliability.
@@
Specifically, we utilize a standard Transformer-Encoder to process the representations of all problem images in an RPM case, alongside $W$ optimizable vectors (where $W$ is a hyperparameter) and a stochastic promoter. Notably, when these $W+1$ vectors are arranged together, they resemble the shape of a glove. The promoter is a vector sampled from a multivariate Gaussian distribution, wherein the mean is optimizable while the variance is fixed at 1.
It is worth noting that we have devised a novel positional encoding embedding approach for the Transformer-Encoder, which serves as the backbone of our RPM generator. This encoding aims to restore the $3\times 3$ image arrangement characteristic of the RPM problem matrix. For the PGM (Progressive Graph Matching) problem, our positional encoding scheme assigns unique encodings to images located on the diagonal and upper triangular regions within the problem matrix of a given PGM case. Correspondingly, the positional encodings for images in the lower triangular region are symmetrical with respect to the diagonal, mirroring the encodings of their counterparts in the upper triangular region. However, it should be acknowledged that this paper does not demonstrate the superiority of D4C on the relatively less challenging RAVEN dataset.
Moving on to the main topic, after obtaining attention results from the Transformer-Encoder, we extract $W$ optimizable vectors from these results and process them into $W$ distributions using a linear layer, as mentioned earlier. Following the extraction of the promoter, it is processed into a probability output using a linear layer and hard-gumbel-softmax activation, yielding highlighted probabilities for the $W$ distributions. Subsequently, we sample representations, referred to as confusing representations, from the highlighted distributions each time. In the technical implementation, we employ the reparameterization trick to draw samples from each distribution, resulting in $W$ sampling outcomes. We then perform a weighted sum of these sampling results using the hard-highlighted probabilities processed by the hard-gumbel-softmax, thus completing the entire mixed-Gaussian sampling process. The methodology outlined in this paper for modeling and training the network is coined as ”D4C glove-train”. A more detailed illustration of the forward process is provided in the accompanying figure referenced as 18.
@@
Finally, in the realm of abstract reasoning problems, particularly those involving generating answer samples, this paper believes that the D4C approach can also make valuable contributions. This opens up new avenues of exploration for future work, presenting challenges that are yet to be tackled.

### v12

@@ VI-A D2C on Bongard-Logo problem
In the Bongard-Logo problem, the determination of image attributes is influenced by the specific contextual concepts associated with each individual case [32]. Distinct concepts require the network to encode specific pixel configurations into different attribute representations, potentially compelling the network to maintain mutually exclusive representation encoding strategies and methods [32]. This scenario often leads to concept confusion within the network, rendering it uncertain about which concept to rely on when processing a Bongard-Logo case [32]. In a nutshell, from the perspective of network, there exists mutual infringement and erroneous coupling among the high-dimensional human concepts in the Bongard-Logo cases. Therefore, this study aims to develop a methodology that effectively decouples the concepts across different cases within the Bongard-Logo problem. We posit that by successfully decoupling such that distinct cases rely on distinct concepts we may alleviate, to a certain extent, the issue of the network losing concept matching when confronted with a Bongard-Logo problem.
Therefore, we strive for mutually exclusive expressions of concepts across all Bongard-Logo cases. By establishing a one-to-one correspondence between problem cases and concepts, we aim to enable the network to avoid uncertainty in determining which concept to rely on when solving a specific case. This approach aims to mitigate the impact of concept confusion on the solution for the Bongard-Logo problem.
@@ VI-A1 D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
The design idea of D2C phase 1 is quite simple, which involves introducing primary group samples from other cases as additional samples into the auxiliary group of the current case. During the implementation process, for each case under consideration, we introduce two distinct cases into it. This process can be regarded as ancestral sampling. This constitutes a plug-and-play approach for training most Bongard-Logo solver such as PMoC[31] and Triple-CFN[32], which is designated as phase 1 of the D2C method in this paper. The process of D2C phase 1 is in Figure 11. As evident from the figure, the ancestral sampling process can be accomplished by performing a misalignment stacking of the sampled Bongard-Logo cases. By incorporating this additional ancestral sampling procedure into the traditional mini-batch training sampling methods, this study constructs a Bongard-logo problem with reorganized concept space for training purposes. This strategy sends a clear and additional signal to Bongard-Logo discriminators that distinct problem cases should rely on differing concepts. This paper aims to utilize this strategy to mitigate concept infringement and erroneous coupling among different Bongard-Logo cases.
@@
It is worth emphasizing that the improvement in model performance provided by D2C is not solely attributed to an increase in the number of negative examples within each contrastive group. This paper posits that the incorporation of negative examples, which assist in delineating concept boundaries among cases, holds the key to enhancing a model’s proficiency in abstract reasoning tasks. Simply adding unhelpful or irrelevant negative examples to the dataset does not significantly improve model performance. In an attempt to enrich the dataset, this study generated a substantial amount of Bongard-logo image data using DDPM [9] and integrated these as auxiliary group sample within the original problems. However, this methodology based on DDPM [9] did not significantly enhance model accuracy.
@@ VI-A2 D2C phase 2: The concept delineation method based on a new loss function term
This paper proposes that, besides changing training data sampling, adding a loss function to limit the similarity of primary group samples’ representations across cases in a Bongard-Logo training batch implicitly fulfills the D2C’s main target.
Specifically, in this paper, we exclusively retain the representation of primary group within the set $\{z_{ij}|\,i\in[1,b],j\in[1,14]\}$, which is $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$. Subsequently, we randomly partition the primary representation of each case within $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$ into two distinct subsets: $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$. By taking the mean of the vectors within $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$, we derive the corresponding vectors $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$. Ultimately, we compute InfoNCE loss function for constraining the similarity among $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$ as the additional loss function item. The specific calculation process can be explained by:
@@
$=-\sum^{b}_{i=1}\log\frac{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}}}{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}+\sum\nolimits_{\tilde{i}=1,\tilde{i}\neq i}^{b}{{e^{({v_{i}}\cdot{v_{\tilde{i}}})/t}}}}}$ (6)
@@
$\text{and}\,\,\{z_{ih}\}\cup\{z_{i\tilde{h}}\}=\{z_{ij}|\,j\in[1,7]\}$
This loss term is to decrease the similarity of primary group representations across different cases. Evidently, D2C employs a ”rather be safe than sorry” approach, which may lead to some overcorrection.
Essentially, D2C’s phase 2 aims to achieve a similar goal as the phase 1 by introducing a new loss function. However, this makes phase 2 mainly compatible with backbones abstractable as $f(z|x)$. This allows most image encoders like ResNet and Triple-CFN [32] to be compatible with full D2C. But PMoC[31], relying on probability measures, is limited to D2C’s phase 1.
@@ VI-B D2C on RPM probelm
Lico-Net can achieve the integration with D2C by introducing images from other RPM cases as incorrect options within the option pool of the current RPM case. It’s worth noting that the application of D2C in RPM discriminators is confined to phase 1 due to the complexity arising in the implementation of phase 2. This process is shown in Figure 12.
@@
VI-C The summary of D2C
It is evident that D2C employs a precautionary approach, aiming to minimize mutual infringement and erroneous coupling among concepts, even at the cost of potential overcorrection. However, we posit that D2C is suitable for addressing the Bongard-Logo problem, where there exists mutual infringement and erroneous coupling of concepts across cases, but it is challenging to apply D2C to the RPM problem, which involves sharing a common set of concepts across cases and focuses on consistent attribute identification.
@@ VII Defense-Driven Discrimination of Distributed Concepts (D4C)
The performances of D2C and D3C on both Bongard-Logo and RPM tasks have stimulated deeper reflections on abstract reasoning challenges. This raises some fascinating questions: Why does modifying negative examples have a profound impact on the solver’s performance specifically in the Bongard-Logo problem? Beyond posing difficulties for the discriminator, what are the additional roles played by positive and negative examples in abstract reasoning challenges? After thoughtful consideration of these aspects, this section introduces the D4C approach.
@@ VII-A Negative samples are not merely intended to confound the discriminator
This paper posits that it is worthwhile to consider the significance of positive and negative examples in abstract reasoning problems. Currently, existing discriminative models for solving RPM and Bongard-Logo problems often adhere to a pattern known as the probability labeling scheme. As the name suggests, when a positive example is submitted to the reasoning problem’s statement, the discriminator will respond with a high probability, whereas the negative examples will be labeled with a low one. We can view this pattern in the following way: the discriminator constructs a solution distribution for the cases presented in abstract reasoning problems. Within this distribution, positive examples are assigned higher probabilities, whereas negative examples are assigned lower ones. The endeavor to train the discriminator can be conceptualized as a process of optimizing and fine-tuning the solution distribution, with the ultimate objective of aligning it more closely with the underlying distribution of correct solutions for the given cases.
To further explain our approach in a clearer and intuitive manner, we assume that the distribution of correct solutions for each case in a reasoning problem follow a Gaussian distribution. Given that some reasoning problems may have multiple correct solutions, it is also possible that the distribution is a mixed Gaussian distribution. Using positive and negative examples to estimate the parameters of a Gaussian distribution through the cross-entropy loss function can be seen as leveraging them to optimize the following function:
@@
Where $x$ represents the positive or negative samples, $\mu$ and $\sigma$ represent the mean and standard deviation of the Gaussian distribution separately. It can be observed that when positive samples are substituted into the formula, we aim for the smallest possible result, which corresponds to minimizing the Euclidean distance between sample $x$ and the mean of the correct solution $\mu$. Conversely, when negative samples are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma$.
Therefore, this paper posits that in reasoning problems, the establishment of positive examples aids deep models in estimating the mean of the correct solution distribution, while negative examples serve to limit the variance. The variance represents the distance between the boundaries of the distribution and its center. Correspondingly, we posit that both positive and negative examples play a crucial role in abstract reasoning tasks. Positive examples assist in determining the centroid of the concepts presented in each case, whereas negative examples contribute to delineating the boundaries of these concepts. Thus, whether negative examples can minimize the variance $\sigma^{2}$ to the size of the correct distribution becomes crucial. If the quality of negative examples is not high enough, the variance cannot be minimized effectively, in other words, it cannot delineate a more precise concept boundary. This underscores the significance of establishing both positive and negative examples in abstract reasoning problems.
This significance reflects a despairing fact that the upper limit of the discriminator’s reasoning accuracy in this abstract reasoning problem is determined by the quality of positive and negative examples in the question.
@@ VII-B High-quality negative samples lead to more rigorous concept boundaries
The aforementioned perspective is the core contribution of this paper. Since negative examples delineate the boundaries of concepts, we posit that enhancing the quality of negative examples equates to delineating more precise concept boundaries. A more precise boundary imply superior discriminator performance. Therefore, this paper aims to device a method that can enhance the quality of negative examples in reasoning problems, and D2C is one such approach.
From the novel perspective on negative samples in reasoning problems, the D2C methodology presented in this study essentially defines a new problem boundary for deep networks. However, we suggest that this approach of defining a new concept boundary, specifically the introduction of images from other cases as negative examples for the current case, can make a significant contribution to problems like Bongard-Logo, where there exists concept infringement and erroneous coupling among different cases. Conversely, for problems like RPM, where different cases share a common concept set, this method may not contribute effectively. In other words, due to the specific characteristics of the Bongard-Logo problem, D2C’s approach has, to some extent, enhanced the quality of negative examples and delineated more precise concept boundaries. However, it cannot replicate these improvements in RPM cases, owing to the distinct problem characteristics of RPM.
In a nutshell, while D2C introduces new negative examples and concept boundaries for reasoning problems, it cannot guarantee the superior quality of these examples or the precision of the new concept boundaries. Hence, this paper proposes the D4C (Defense-Driven Discrimination of Distributed Concepts) method, which deliberately seeks out higher-quality negative examples for reasoning problems, thereby defining more accurate concept boundaries. The underlying idea of D4C is to generate higher-quality negative samples for abstract reasoning problems, aiming to delineate a more appropriate problem boundary for the discriminator compared to the original problem setup in abstract reasoning. By leveraging novel mechanisms and strategies, D4C strives to deliver more precise and reliable definitions of concept boundaries, fostering further advancements in the realm of abstract reasoning.
@@ VII-C Impelmention of D4C method
Effective Bongard-Logo discriminators abound, with our preceding works, such as PMoC[31] and Triple-CFN[32], achieving notable successes. Regarding discriminators capable of handling adversarial tasks, the discriminator must undertake the task of measuring the distance between samples generated by the generator and real samples. This necessitates that the discriminator satisfies Lipschitz continuity to a certain extent[36]. Therefore, in this paper, we choose the spectrum-normalized PMoC and Triple-CFN as discriminators for our D4C, and design generators to refine them.
@@
Specifically, we employ a standard Transformer encoder to encode a Gaussian distribution based on the representations of the primary samples in a Bongard-Logo case, along with an optimizable vector. This encoding approach bears similarity to D3C, wherein we extract the optimizable vector from the attention results and process it through a linear feed-forward layer to obtain the mean and logarithm of variance for a Gaussian distribution. Subsequently, we sample multiple representations from this distribution as confusing representations to challenge the discriminator, which must strive to counter them. The generator’s training objective is to generate representations for Bongard-Logo cases that achieve higher scores in the discriminator compared to the representations of the original negative examples in the case. Conversely, the discriminator’s training objective is unambiguous: to classify these generated representations as negative examples. The aforementioned approach can be seamlessly applied to the PMoC model, providing a straightforward and effective means of enhancing its performance. The integration method between the generator and PMoC is illustrated in Figure 13.
@@
For the generator in Triple-CFN[32], it is essential to decay the generated representations. Specifically, when the generator with the aforementioned structure produces a confusing representation for Triple-CFN, this confusing representation must be weighted and summed with the representations in the auxiliary group. After taking the average of the representations in the auxiliary group, it is weighted and summed with the confusing representation in a ratio of $0.9:0.1$ to decay the confusing representation. Due to the mechanism of Triple-CFN, this decay process is deemed necessary, which defines an image’s reasoning identity based on the direction of its representation. This mechanism makes it easier for a Transformer Encoder-based generator to generate realistic confusing representations, potentially causing Triple-CFN to collapse. Therefore, we have designed the aforementioned decay process. The integration method between the generator and Triple-CFN is illustrated in Figure 14.
@@
Lico-Net will encounter challenges posed by the generator, which produces confusing representations aimed at completing the $3\times 3$ reasoning matrix. To address these challenges, Lico-Net will undertake an assessment of the internal consistency within the progressive patterns of the $3\times 3$ reasoning matrix, which are completed by these confusing representations. Any inconsistencies or irregularities identified within these representations will be flagged as negative samples, indicating incorrect responses. Through this rigorous process, Lico-Net strives to accurately identify and reject confusing representations, thereby augmenting its overall precision and reliability.
@@
Specifically, we utilize a standard Transformer-Encoder to process the representations of all images in the statement of an RPM case and $N$ pattern vectors from the statement encoded by Lico-Net, alongside $W$ optimizable vectors (where $W$ is a hyperparameter) and a stochastic promoter. The promoter is a vector sampled from a multivariate Gaussian distribution, wherein the mean is optimizable while the variance is fixed at 1. The Transformer-Encoder needs to process $N+W+1+8$ vectors in total. Notably, when these $W$ vectors and the promoter are arranged together, they resemble the shape of a glove.
It is worth noting that we have devised a novel positional encoding embedding approach for the Transformer-Encoder, which serves as the backbone of our RPM generator. This encoding aims to restore the $3\times 3$ image arrangement characteristic of the RPM problem matrix. For the PGM problem, our positional encoding scheme assigns unique encodings to images located on the diagonal and upper triangular regions within the problem matrix of a given PGM case. Correspondingly, the positional encodings for images in the lower triangular region are symmetrical with respect to the diagonal, mirroring the encodings of their counterparts in the upper triangular region. However, it should be acknowledged that this paper does not demonstrate the superiority of D4C on the relatively less challenging RAVEN dataset.
Moving on to the main topic, after obtaining attention results from the Transformer-Encoder, we extract $W$ optimizable vectors from these results and process them into $W$ distributions using a linear layer, like D3C method. Following the extraction of the promoter, it is processed into a probability output using a linear layer and hard Gumbel-Softmax activation, yielding highlighted probabilities for the $W$ distributions. Subsequently, we sample representations, referred to as confusing representations, from the highlighted distributions each time. In the technical implementation, we employ the reparameterization trick to draw samples from each distribution, resulting in $W$ sampling outcomes. We then perform a weighted sum of these sampling results using the hard-highlighted probabilities processed by the hard Gumbel-Softmax, thus completing the entire mixed-Gaussian sampling process. The methodology outlined in this section is coined as “D4C glove-train”. The forward process of D4C on Lico-Net is shown in the figure 15. In figure 15, “pattern extractor” is illustrated in figure 6 and the process of “pattern scoring” is shown in figure 7.
@@
Overall, compared to traditional methods, D4C is a fundamentally different approach. Previous studies on solving abstract reasoning problems[23, 21, 28, 32] have primarily focused on constructing solvers that structurally align with the inductive bias in the problems, including our Lico-Net. However, our D4C reconsiders the significance of negative examples in reasoning problems, and designs a solution based on this introspection. As a bonus, the glove-shaped generator in D4C provides suggestions for simulating the random sampling process of a mixed Gaussian distribution.